a child is sitting on a swing .its minimum and maximum heights from the ground are 0.75m and 2m resp. its maximum speed will be
1
Expert's answer
2012-12-19T10:54:19-0500
T - kinetic energy
U - potential energy
The Law of Conservation of Energy:
T+U = const
U = mgh
T = (mv^2)/2
T_max+U_min = 0 + U_max
(m*(v_max)^2)/2 = mg(h_max - h_min)
v_max = sqrt(2g(h_max - h_min) = sqrt(2*10*1.25) = sqrt(25) = 5 m/s

Answer : v_max = 5 m/s